OpenStudy (anonymous):

Find the work done in breaking a water drop of radius 1 mm into 1000 drops. Given the surface tension of water is 72 × 10-3 N/m?

OpenStudy (anonymous):

first find radius of each of 1000 drops as volume of 1 drop(first)=volume of 1000 single drops 4/3pi*Rcube=4/3*pi*rcube*1000 work is done by surface energy=surface tension*increse in surface area surface area of 1000 new drops is definitely greater than surface area of a single water drop surface area=4*pi*rsquare